# Efficient Diffusion Models: A Survey

### v1

@@ 3.1.3 Data-Dependent Adaptive Priors
To accelerate the generation process in diffusion models and improve the quality of generated samples, better initialization of priors tailored to specific tasks and datasets can be utilized. This approach leverages data-dependent adaptive priors, making the generation process more efficient and aligning the generated samples more closely with true data distribution. Recent studies have explored how data-dependent adaptive priors can enhance diffusion models.
@@
As illustrated in Figure 4, data-dependent adaptive priors can be applied across different modalities, such as speech, graph, and trajectory, to enhance the diffusion process. By aligning priors with data distributions specific to each modality, the model can generate outputs that are better tailored to the underlying structure of the data. In traditional diffusion models, the prior is typically assumed to be a standard Gaussian distribution $p(z)=\mathcal{N}(0,I)$. However, this may not always align well with the actual data distribution, potentially introducing inefficiencies. By leveraging data-dependent adaptive priors based on $X$, where $X$ represents the data, the model can achieve better initialization and faster convergence, without relying solely on a standard Gaussian assumption. Lee et al. (2021) introduce PriorGrad, which improves diffusion models for speech synthesis by using an adaptive prior derived from data statistics based on conditional information. This method increases the efficiency of the denoising process, leading to faster convergence and inference speed, while also improving perceptual quality and robustness with smaller network capacities.
Digress (Vignac et al., 2022) presents a discrete denoising diffusion model focused on graph generation. This model leverages data-dependent priors to better capture the discrete nature of graph data, significantly enhancing the quality of generated graphs, particularly for applications such as chemical molecular structures and social networks. DecompDiff (Guan et al., 2023) introduce decomposed priors, modeling different structural components of drug molecules separately to improve diffusion models. This approach enhances the accuracy of generating viable drug candidates by better capturing molecular structure information. As illustrated in Figure 5, Mao et al. (2023) propose the Leapfrog Diffusion Model for stochastic trajectory prediction, introducing a leapfrog initializer that uses adaptive priors to skip multiple denoising steps. This method significantly accelerates inference while maintaining high prediction accuracy, making it useful for real-time applications such as autonomous driving and robotic navigation.
@@
3.1.4 Rectified Flow
Figure 6: Illustration of the rectified flow.
As illustrated in Figure 6, Rectified Flow, proposed by (Liu et al., 2022b; Liu, 2022), introduces a method for training ordinary differential equation (ODE) models by learning straight transport paths between two distributions, $\pi_{0}$ and $\pi_{1}$. The key idea is to minimize the transport cost by ensuring that the learned trajectory between these two distributions follows the most direct route, which can be computationally efficient to simulate. Unlike traditional diffusion models, which may follow roundabout paths, Rectified Flow leverages a simpler optimization problem to create a straight flow, thereby improving both training efficiency and the quality of the generated samples.
Building upon this foundation, InstaFlow (Liu et al., 2023b) apply the Rectified Flow concept to text-to-image generation, achieving a significant breakthrough. InstaFlow represents a major advancement in efficient diffusion models, which are capable of high-quality image generation in just one step. It applied Rectified Flow to large-scale datasets and complex models like Stable Diffusion, introduced a novel text-conditioned pipeline for one-step image generation, and achieved an FID score of 23.3 on MS COCO 2017-5k. InstaFlow’s success highlights the potential of Rectified Flow in dramatically reducing the computational cost of diffusion models while maintaining high output quality.
@@ 3.2 Efficient Fine-Tuning
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw=hidden-draw, rounded corners, align=left, text centered, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=10em,font=,, where level=2text width=30em,font=,, where level=3text width=30em,font=,, [ Efficient Fine-tuning [ LoRA, fill=blue!10 [ LoRA (Hu et al., 2021), LCM-LoRA (Luo et al., 2023b),
Concept Sliders (Gandikota et al., 2023), LoRA-Composer
 Yang et al. (2024), LoRA conditioning (Choi et al., 2024b), ] ] [ Adapter, fill=blue!10 [ T2I-Adapter (Mou et al., 2024), IP-Adapter (Ye et al., 2023),
ParaTAA (Tang et al., 2024), CTRL-Adapter
 (Lin et al., 2024a), SimDA (Xing et al., 2024), ] ] [ ControlNet, fill=blue!10 [ ControlNet (Zhang et al., 2023c), ControlNet++ (Li et al., 2025),
ControlNet-XS (Zavadski et al., 2023), ControlNeXt
 (Peng et al., 2024a), Uni-ControlNet (Zhao et al., 2024),
UniControl (Qin et al., 2023), ] ] ]
@@
Fine-tuning pre-trained diffusion models can be computationally expensive. To address this, we categorize efficient fine-tuning techniques into LoRA, Adapters, and ControlNet, based on their mechanisms for reducing resource consumption. This classification reflects differences in how these methods optimize parameters, enable task-specific adaptation, and incorporate spatial conditioning signals, as summarized in Figure 7.
@@ 3.2.1 LoRA
The complete process is illustrated in Figure 8. A key advantage of this design lies in its deployment efficiency, where the explicit computation and storage of $W=W_{0}+BA$ enables standard inference procedures without introducing additional latency. Originally proposed for fine-tuning Large Language Models (LLMs), LoRA has demonstrated remarkable parameter efficiency and memory reduction in model adaptation.
While predominantly utilized in LLM fine-tuning, recent research has extended its application to diffusion models, indicating its potential as a versatile adaptation technique across different deep learning architectures. LCM-LoRA (Luo et al., 2023b) proposes a universal acceleration approach for diffusion models. As shown in Figure 9, this method achieves fast sampling by adding an Acceleration vector $\tau_{LCM}$ to the Base LDM Rombach et al. (2022). This module adopts LoRA (Hu et al., 2021) to attach low-rank matrices to the original model without architectural modifications. For customized diffusion models that are fine-tuned for specific text-to-image generation tasks, the task-specific LoRA ($\tau^{\prime}$) and acceleration LoRA ($\tau_{LCM}$) can be linearly combined through Eq.(18) to achieve fast inference while maintaining generation quality. More importantly, it provides a plug-and-play solution that reduces sampling steps from dozens to around 4, while maintaining compatibility with any pre-trained text-to-image diffusion model.
@@
where $\epsilon_{\theta}$ represents the denoising model’s prediction, and $\eta$ is a guidance scale. With this formulation, the method enables smooth control over concept strength through the guidance scale $\eta$ while maintaining concept independence in the learned directions. By leveraging LoRA’s parameter-efficient nature, it achieves precise attribute manipulation with minimal computational overhead.
@@ 3.2.2 Adapter
Adapters are lightweight modules designed to enable efficient task adaptation by introducing small network layers into pre-trained models, allowing task-specific feature learning while keeping the original weights frozen. As illustrated in Figure 10, adapter layers are placed within the transformer block, positioned between normalization and feed-forward layers. Each adapter module consists of a down-projection, nonlinearity, and up-projection, which generates task-specific transformations without altering the core model’s structure. This design significantly reduces memory and computational requirements, making adapters well-suited for tasks requiring lightweight parameter updates, such as text-to-image generation (T2I) and simulated domain adaptation (SimDA).
IP-Adapter (Ye et al., 2023) enhances the consistency and visual quality of text-to-image generation by incorporating image prompts. Unlike T2I-Adapter (Mou et al., 2024), which provides structural guidance through sketches or depth maps, IP-Adapter focuses on capturing visual details from an input image, making it ideal for tasks requiring high visual consistency with a reference image. This adapter processes the input image prompt into latent features, allowing the generation model to capture visual information from the target image and maintain detail alignment throughout the generation process. In its workflow, the image prompt is first mapped into the latent space and then processed through convolution and normalization modules within the adapter, enabling the model to utilize these features during inference. This setup enables the generation model to draw rich visual information from the image prompt, making IP-Adapter particularly suitable for tasks requiring high detail consistency, such as generating images with a style similar to the input image.
@@
CTRL-Adapter (Lin et al., 2024a) is designed to enhance attribute control during generation by guiding specific attributes such as emotion or object type, enabling precise customization in generated results. Unlike T2I-Adapter (Mou et al., 2024) and IP-Adapter (Ye et al., 2023), which focus on structural and detail consistency respectively, CTRL-Adapter is tailored to provide diversity control for the generation model. For example, as illustrated in Figure 11, the IP-Adapter architecture employs a decoupled cross-attention strategy, where only newly added modules are trained while the pre-trained text-to-image model remains frozen. In contrast, CTRL-Adapter can adjust the style of generated images based on specified emotions or object types, achieving controllable content generation without altering the core architecture of the model. This makes CTRL-Adapter particularly suitable for tasks requiring high customization in generation, such as emotion-driven text generation or stylized image synthesis.
@@ 3.2.3 ControlNet
ControlNet (Zhang et al., 2023c) and its derivatives represent a significant advancement in adding spatial conditioning controls to pre-trained text-to-image diffusion models. The original ControlNet architecture (Zhang et al., 2023c), as illustrated in Figure 12, presents a novel approach to integrating various spatial conditions—such as scribbles, edge maps, open-pose skeletons, or depth maps—into the generative process while preserving the robust features of pre-trained diffusion models. The architecture employs zero convolution layers that gradually develop parameters without disrupting the pre-trained model’s stability. This design enables versatile conditioning, allowing the model to effectively leverage different types of spatial information. Through these conditioning methods, ControlNet demonstrates a remarkable ability to guide generation with fine-grained control over structure, style, and composition. Building upon this foundation, several works have proposed improvements and alternatives. ControlNet++ (Li et al., 2025) addresses the challenge of alignment between generated images and conditional controls by introducing pixel-level cycle consistency optimization. Through a pre-trained discriminative reward model and an efficient reward strategy involving single-step denoised images, it achieves significant improvements in control accuracy, with notable gains in metrics such as mIoU (11.1%), SSIM (13.4%), and RMSE (7.6%) across various conditioning types. ControlNet-XS (Zavadski et al., 2023) reimagines the control system by enhancing the communication bandwidth between the controlling network and the generation process. This redesign not only improves image quality and control fidelity but also significantly reduces the parameter count, resulting in approximately twice the speed during both inference and training while maintaining competitive performance in pixel-level guidance tasks. The field has also seen efforts to unify multiple control capabilities. UniControl (Qin et al., 2023) introduces a task-aware HyperNet approach that enables a single model to handle diverse visual conditions simultaneously. Similarly, Uni-ControlNet (Zhao et al., 2024) proposes a unified framework supporting both local controls and global controls through just two additional adapters, significantly reducing training costs and model size while maintaining high performance.
Figure 12: Illustration of ControlNet.
Most recently, ControlNeXt (Peng et al., 2024a) has pushed the boundaries of efficiency even further by introducing a streamlined architecture that minimizes computational overhead. It replaces the traditional heavy additional branches with a more concise structure and introduces Cross Normalization (CN) as an alternative to zero convolutions. This approach achieves fast and stable training convergence while reducing learnable parameters by up to 90% compared to previous methods.

### v2

@@ 3.1.3 Training Tricks
Data-Dependent Adaptive Priors. To enhance the training efficiency of diffusion models and improve the quality of generated samples, data-dependent adaptive priors can be utilized to tailor the prior distribution to specific tasks and datasets. This approach leverages priors that better align with the data distribution, thereby accelerating the training process and ensuring that generated samples more closely match the true data distribution. Recent studies have explored how data-dependent adaptive priors can improve the training of diffusion models.
@@
As a method under efficient training (see Figure 5), data-dependent adaptive priors can be applied across various modalities, such as speech, graphs, and trajectories. By aligning the prior with the data distribution specific to each modality, the model can achieve faster convergence during training while producing outputs that better reflect the underlying data structure. In traditional diffusion models, the prior is typically assumed to be a standard Gaussian distribution $p(z)=\mathcal{N}(0,I)$. However, this assumption may not align well with the actual data distribution, potentially leading to inefficiencies in training. By constructing data-dependent adaptive priors based on the data $X$, the model can achieve better initialization during training, accelerating convergence without relying solely on the standard Gaussian assumption.
Lee et al. (2021) introduced PriorGrad, which enhances the training of diffusion models for speech synthesis by using an adaptive prior derived from conditional data statistics. This method significantly improves the efficiency of the denoising process during training, leading to faster convergence while enhancing the perceptual quality and robustness of generated samples, even with smaller network capacities.
Vignac et al. (2022) proposed DiGress, a discrete denoising diffusion model for graph generation. By leveraging data-dependent priors, this model better captures the discrete nature of graph data, improving training efficiency and the quality of generated graphs, particularly for applications like chemical molecular structures and social networks.
In drug design, Guan et al. (2023) introduced DecompDiff, which improves the training of diffusion models by using decomposed priors to separately model different structural components of drug molecules. This approach enhances the model’s ability to capture molecular structure information during training, leading to the generation of more accurate drug candidates.
As shown in Figure 6Mao et al. (2023) proposed the Leapfrog Diffusion Model for stochastic trajectory prediction, introducing a leapfrog initializer based on adaptive priors to skip multiple denoising steps, accelerating training while maintaining accuracy for real-time applications like autonomous driving. Building on this, Fu et al. (2025) developed MoFlow, a one-step flow matching model with IMLE-based distillation for human trajectory forecasting. MoFlow employs a novel flow matching loss to ensure accuracy and diversity in predicted trajectories, using data-dependent adaptive priors based on past trajectories and interactions to enhance alignment with the data distribution. Its IMLE distillation achieves a 100x faster one-step student model with comparable performance. Similarly, Jiang et al. (2024) introduced SceneDiffuser, a scene-level diffusion model using amortized diffusion to optimize efficiency for driving simulation, supporting both initialization and rollout. These advances highlight the role of adaptive priors in boosting training efficiency for real-time multimodal applications.
@@
Extra Loss Supervision. To further enhance the training efficiency and sample quality of diffusion models, incorporating additional loss supervision has proven to be an effective strategy. Beyond standard denoising objectives, extra loss terms can guide the model toward better alignment with data distributions, accelerate convergence, and improve the robustness of generated outputs. Recent advancements have demonstrated the power of tailoring loss functions to leverage external knowledge or address specific optimization challenges.
One notable approach introduces REPresentation Alignment (REPA) (Yu et al., 2024a), a regularization technique that aligns the hidden states of the denoising network with representations from a pre-trained visual encoder, such as DINOv2. By adding a loss term that enforces this alignment, REPA ensures that the diffusion Transformer (e.g., DiT or SiT) leverages high-quality external visual priors rather than learning representations from scratch. This supervision accelerates training—achieving over 17.5$\times$ speedup in some cases—and boosts generation quality, reaching a state-of-the-art FID of 1.42 on ImageNet 256x256 with classifier-free guidance. The additional loss acts as a bridge between the model’s internal learning and established visual knowledge, making training both faster and more effective.
Similarly, another method tackles the optimization dilemma in latent diffusion models (LDMs) by introducing extra loss supervision through the VA-VAE framework (Yao et al., 2025). This approach aligns the latent space of a variational autoencoder with a pre-trained vision foundation model, using a loss term to ensure the latent representations capture the data distribution more effectively. This supervision mitigates the trade-off between reconstruction fidelity and generation quality, a common challenge in LDMs where increasing latent dimensionality often hampers generative performance. Combined with an optimized LightningDiT architecture, this method achieves an FID of 1.35 on ImageNet 256x256, with training convergence accelerated by over 21$\times$ compared to baseline DiT models. The extra loss supervision in VA-VAE enhances the model’s ability to balance reconstruction and generation tasks, leading to both efficiency gains and superior sample quality.
Noise Schedule. Noise schedule is a crucial component of diffusion models, governing how noise is added during the forward process and removed during the reverse process. Optimizing the noise schedule can significantly enhance the training efficiency of diffusion models by enabling faster convergence and more effective learning of data distributions. Denoising Diffusion Probabilistic Models (DDPM) (Ho et al., 2020) introduced a linear noise schedule that gradually decreases the variance of the noise added in the forward process, as defined in Eq.(1). However, the linear schedule requires calculating complex noise terms across numerous timesteps, which can slow down the training process and highlight the need for more efficient noise schedule designs. As shown in Figure 7, efficient noise schedules can be classified into two main categories: systematic noise addition and dynamic noise adjustment.
Figure 7: Illustration of two categories of noise schedules.
One approach involves systematically adding noise during the training process at predefined intervals or according to specific levels. DDPM (Ho et al., 2020) employs a linear noise schedule, where the noise variance changes deterministically over time, serving as a foundational example. Building on this, the Improved Denoising Diffusion Probabilistic Model (IDDPM) (Nichol & Dhariwal, 2021) introduces a cosine noise schedule, defined as
$\beta_{t}=1-\frac{\cos\left(\frac{t/T+s}{1+s}\cdot\frac{\pi}{2}\right)}{\cos\left(\frac{s}{1+s}\cdot\frac{\pi}{2}\right)},$ (24)
where $t$ is the current timestep, $T$ is the total number of timesteps, and $s$ is a small positive constant for smoothing initial noise addition. The cosine noise schedule optimizes the noise distribution during training, enabling the model to learn data structures more effectively and achieve faster convergence. However, it allocates computational resources evenly across all noise intensities, which may not prioritize the most critical regions for training.
To address this inefficiency, Hang & Gu (2024) proposed the Laplace noise schedule, which enhances training efficiency by increasing the sampling frequency around critical noise regions. This approach ensures that the model focuses computational resources on medium noise intensities, which are more effective for learning data structures and removing noise, leading to faster convergence and improved sample quality during training. The Laplace schedule balances noise addition across timesteps, resulting in a more robust training process.
For text generation, traditional diffusion models often add uniform Gaussian noise to each word, which fails to leverage linguistic features and increases computational burden during training. To address this, Chen et al. (2023a) introduced a soft-masking noise strategy that gradually adds noise based on word importance, measured by term frequency and information entropy. Using a square-root noise schedule (Li et al., 2022), this method incrementally increases noise levels, stabilizing the training process and enabling the model to learn linguistic structures more effectively from the initial latent variable $X_{0}$ to noisy variables $X_{1:T}$.
In contrast, another set of methods dynamically adjusts the noise schedule based on the model’s state or data during training. Song et al. (2020a) developed Denoising Diffusion Implicit Models (DDIM), which improve the noise schedule in DDPM by introducing a non-Markovian forward process, defined as
$x_{t-1}=\sqrt{\alpha_{t-1}}\left(x_{t}-\sqrt{1-\alpha_{t}}\epsilon_{\theta}(x_{t})\right)+\sqrt{1-\alpha_{t-1}-\sigma_{t}^{2}}\epsilon_{\theta}(x_{t})+\sigma_{t}\epsilon_{t},$ (25)
where $\alpha_{t}$ controls the noise addition over time, and $\sigma_{t}$ is dynamically adjusted based on the current state $x_{t}$ and initial state $x_{0}$. This dynamic adjustment leverages the entire trajectory, enabling more precise noise control during training, which enhances the model’s ability to learn data distributions efficiently and improves training convergence.
Inspired by DDIM, Yue et al. (2024) proposed ResShift, which introduces a noise schedule that constructs a shorter Markov chain by shifting residuals between high-resolution (HR) and low-resolution (LR) images instead of adding Gaussian noise. The noise schedule is defined as
$\sqrt{\eta_{t}}=\sqrt{\eta_{1}}\times b_{0}^{\beta_{t}},\quad t=2,\ldots,T-1,$ (26)
where
$\beta_{t}=\left(\frac{t-1}{T-1}\right)^{p}\times(T-1),\quad b_{0}=\exp\left(\frac{1}{2(T-1)}\log\frac{\eta_{T}}{\eta_{1}}\right),$ (27)
with $T$ as the total timesteps, $t$ as the current timestep, $p$ controlling the growth rate of $\sqrt{\eta_{t}}$, and $\eta_{1}$ and $\eta_{T}$ as the initial and final noise levels. This non-uniform noise progression allows the model to focus training on key regions, improving convergence and learning efficiency for super-resolution tasks.
To further optimize training, Li et al. (2024d) proposed Immiscible Diffusion, inspired by the physical phenomenon of immiscibility. Unlike traditional methods that diffuse each image across the entire noise space, this approach reassigns noise to images within a mini-batch to minimize the distance between image-noise pairs. By matching each image with nearby noise, Immiscible Diffusion reduces the complexity of denoising during training, enabling the model to learn data structures more efficiently.
For text generation, traditional methods (Gong et al., 2022) often use fixed noise schedules, requiring recalculation of the input sequence at each timestep, which leads to inefficient training. In contrast, Yuan et al. (2024) introduced a dynamic noise adjustment technique that records loss values at each timestep during training and uses linear interpolation to map these losses to noise schedule parameters. This adaptive approach optimizes noise levels at each timestep and token position, improving training efficiency and enabling the model to better capture linguistic features.
@@ 3.2 Efficient Fine-Tuning
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw=hidden-draw, rounded corners, align=left, text centered, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=8em,font=,, where level=2text width=30em,font=,, where level=3text width=30em,font=,, [ Efficient Fine-tuning [ LoRA, fill=blue!10 [ LoRA (Hu et al., 2021), Concept Sliders (Gandikota et al., 2023),
LCM-LoRA (Luo et al., 2023b), LoRA-Composer Yang et al. (2024),
LoRA conditioning (Choi et al., 2024b), leaf, text width=32em ] ] [ Adapter, fill=blue!10 [ T2I-Adapter (Mou et al., 2024), IP-Adapter (Ye et al., 2023),
ParaTAA (Tang et al., 2024), CTRL-Adapter (Lin et al., 2024a),
SimDA (Xing et al., 2024), leaf, text width=32em ] ] [ ControlNet, fill=blue!10 [ ControlNet (Zhang et al., 2023d), ControlNet++ (Li et al., 2025),
ControlNet-XS (Zavadski et al., 2023), ControlNeXt (Peng et al., 2024a),
Uni-ControlNet (Zhao et al., 2024), UniControl (Qin et al., 2023), leaf, text width=32em ] ] ]
@@
Fine-tuning pre-trained diffusion models demands efficient resource use, leading to distinct strategies: LoRA, Adapters, and ControlNet. LoRA stands out for its low-rank parameter updates, slashing memory needs by up to 90% and enabling rapid inference, though it struggles with nuanced spatial control. Adapters, by contrast, inject lightweight modules for task-specific tweaks, excelling in conditional guidance (e.g., sketches via T2I-Adapter) but relying on input quality, which can limit robustness. ControlNet, however, leverages additional network branches for precise spatial conditioning, offering unmatched control over structure and style, yet at the cost of higher computational load—mitigated in variants like ControlNet-XS. These trade-offs are explored in detail in the subsections.
@@ 3.2.1 LoRA
The complete process is illustrated in Figure 8. A key advantage of this design lies in its deployment efficiency, where the explicit computation and storage of $W=W_{0}+BA$ enables standard inference procedures without introducing additional latency. Originally proposed for fine-tuning Large Language Models (LLMs), LoRA has demonstrated remarkable parameter efficiency and memory reduction in model adaptation. While predominantly utilized in LLM fine-tuning, recent research has extended its application to diffusion models, indicating its potential as a versatile adaptation technique across different deep learning architectures. LCM-LoRA (Luo et al., 2023b) proposes a universal acceleration approach for diffusion models. As shown in Figure 10, this method achieves fast sampling by adding an Acceleration vector $\tau_{LCM}$ to the Base LDM Rombach et al. (2022). This module adopts LoRA (Hu et al., 2021) to attach low-rank matrices to the original model without architectural modifications. For customized diffusion models that are fine-tuned for specific text-to-image generation tasks, the task-specific LoRA ($\tau^{\prime}$) and acceleration LoRA ($\tau_{LCM}$) can be linearly combined through Eq.(30) to achieve fast inference while maintaining generation quality. More importantly, it provides a plug-and-play solution that reduces sampling steps from dozens to around 4, while maintaining compatibility with any pre-trained text-to-image diffusion model.
@@
Here, $\epsilon_{\theta}$ represents the denoising model’s prediction, and $\eta$ is a guidance scale. With this formulation, the method enables smooth control over concept strength through the guidance scale $\eta$ while maintaining concept independence in the learned directions. By leveraging LoRA’s parameter-efficient nature, it achieves precise attribute manipulation with minimal computational overhead.
@@ 3.2.2 Adapter
Adapters are lightweight modules designed to enable efficient task adaptation by introducing small network layers into pre-trained models, allowing task-specific feature learning while keeping the original weights frozen. As illustrated in Figure 11, adapter layers are placed within the transformer block, positioned between normalization and feed-forward layers. Each adapter module consists of a down-projection, nonlinearity, and up-projection, which generates task-specific transformations without altering the core model’s structure.
This design significantly reduces memory and computational requirements, making adapters well-suited for tasks requiring lightweight parameter updates, such as text-to-image generation (T2I) and simulated domain adaptation (SimDA).
@@
IP-Adapter (Ye et al., 2023) enhances the consistency and visual quality of text-to-image generation by incorporating image prompts. Unlike T2I-Adapter (Mou et al., 2024), which provides structural guidance through sketches or depth maps, IP-Adapter focuses on capturing visual details from an input image, making it ideal for tasks requiring high visual consistency with a reference image. This adapter processes the input image prompt into latent features, allowing the generation model to capture visual information from the target image and maintain detail alignment throughout the generation process. In its workflow, the image prompt is first mapped into the latent space and then processed through convolution and normalization modules within the adapter, enabling the model to utilize these features during inference. This setup enables the generation model to draw rich visual information from the image prompt, making IP-Adapter particularly suitable for tasks requiring high detail consistency, such as generating images with a style similar to the input image. CTRL-Adapter (Lin et al., 2024a) is designed to enhance attribute control during generation by guiding specific attributes such as emotion or object type, enabling precise customization in generated results. Unlike T2I-Adapter (Mou et al., 2024) and IP-Adapter (Ye et al., 2023), which focus on structural and detail consistency respectively, CTRL-Adapter is tailored to provide diversity control for the generation model.
@@
For example, as illustrated in Figure 12, the IP-Adapter architecture employs a decoupled cross-attention strategy, where only newly added modules are trained while the pre-trained text-to-image model remains frozen. In contrast, CTRL-Adapter can adjust the style of generated images based on specified emotions or object types, achieving controllable content generation without altering the core architecture of the model. This makes CTRL-Adapter particularly suitable for tasks requiring high customization in generation, such as emotion-driven text generation or stylized image synthesis.
@@ 3.2.3 ControlNet
ControlNet (Zhang et al., 2023d) and its derivatives represent a significant advancement in adding spatial conditioning controls to pre-trained text-to-image diffusion models. The original ControlNet architecture (Zhang et al., 2023d), as illustrated in Figure 13, presents a novel approach to integrating various spatial conditions—such as scribbles, edge maps, open-pose skeletons, or depth maps—into the generative process while preserving the robust features of pre-trained diffusion models. The architecture employs zero convolution layers that gradually develop parameters without disrupting the pre-trained model’s stability. This design enables versatile conditioning, allowing the model to effectively leverage different types of spatial information. Through these conditioning methods, ControlNet demonstrates a remarkable ability to guide generation with fine-grained control over structure, style, and composition. Building upon this foundation, several works have proposed improvements and alternatives. ControlNet++ (Li et al., 2025) addresses the challenge of alignment between generated images and conditional controls by introducing pixel-level cycle consistency optimization. Through a pre-trained discriminative reward model and an efficient reward strategy involving single-step denoised images, it achieves significant improvements in control accuracy, with notable gains in metrics such as mIoU (11.1%), SSIM (13.4%), and RMSE (7.6%) across various conditioning types. ControlNet-XS (Zavadski et al., 2023) reimagines the control system by enhancing the communication bandwidth between the controlling network and the generation process. This redesign not only improves image quality and control fidelity but also significantly reduces the parameter count, resulting in approximately twice the speed during both inference and training while maintaining competitive performance in pixel-level guidance tasks. The field has also seen efforts to unify multiple control capabilities. UniControl (Qin et al., 2023) introduces a task-aware HyperNet approach that enables a single model to handle diverse visual conditions simultaneously. Similarly, Uni-ControlNet (Zhao et al., 2024) proposes a unified framework supporting both local controls and global controls through just two additional adapters, significantly reducing training costs and model size while maintaining high performance. Most recently, ControlNeXt (Peng et al., 2024a) has pushed the boundaries of efficiency even further by introducing a streamlined architecture that minimizes computational overhead. It replaces the traditional heavy additional branches with a more concise structure and introduces Cross Normalization (CN) as an alternative to zero convolutions. This approach achieves fast and stable training convergence while reducing learnable parameters by up to 90% compared to previous methods.